please help me to solve
2x-y=4
2x-y=3 by substitution

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

There's no solution, the 2 eqns are inconsistent.
Both say 2x-y=
1 says 4
2 says 3
----------
If you graph them, they're parallel lines, no intersection, no solution.

Answer by richwmiller(17219)

when we solve by substitution we solve one of the equations for x or y and then substitute that in the other equation
so if we solve 2x-y=4 for y, we will put the result in 2x-y=3 for y.
let's solve for y
2x-y=4
add y to both sides
2x=4+y
subtract 4 from both sides
2x-4=y
now substitute 2x-4 for y in 2x-1y=3
2x-1(2x-4)=3
Be careful of the signs
minus a plus is minus
but minus a minus is plus
if there is no sign it is a plus